OpenStudy (anonymous):

integrate (((2/3)-(2x/3))/(x^2+x+1))

OpenStudy (anonymous):

just sub divide then cross multiply and add the outliers to x.

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \frac{ 2 }{ 3 }-\frac{ 2x }{ 3 }}{ x^{2}+x+1}\\]

OpenStudy (anonymous):

no quite

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

first you need to factories quadratically

OpenStudy (anonymous):

and use calculus to find the sectum

OpenStudy (anonymous):

factorise quadratically?

OpenStudy (anonymous):

find sectum?

OpenStudy (anonymous):

yes.. u subdivide to x then factorise the quadratic equation on the right.

OpenStudy (anonymous):

split them into 2 different equations you have [(2/3) / (x^2+x+1)] - [2x/3 / (x^2+x+1)] simplify [2x^2 + 2x + 1]/3 - [2x^3 + 2x^2 + 2x] / 3 simplify more 2/3x^2 + 2/3x + 1/3 - 2/3x^3 - 2/3x^2 - 2/3x and your integrating that which is 2/9x^3 + 1/3x^2 + 1/3x - 1/6x^4 - 2/9x^3 -1/3x^2 + c simplify -x^4/6 + x/3 + C is your answer

OpenStudy (anonymous):

then you find the sectum

OpenStudy (anonymous):

and cross multiply.. understood

OpenStudy (anonymous):

? 0

OpenStudy (anonymous):

btw jayz you made a huge mistake

OpenStudy (anonymous):

you forgot to add x to the 3xsquared when you were subdividing..

OpenStudy (sirm3d):

\[\large \frac{ \frac{ 2 }{ 3 }-\frac{ 2x }{ 3 } }{ x^2+x+1 }=-\frac{ 1 }{ 3 }\frac{ 2x-2 }{ x^2+x+1 }\]\[\large = -\frac{ 1 }{ 3 }\frac{ 2x+1-3 }{ x^2+x+1 }=-\frac{ 1 }{ 3 }\left( \frac{ 2x+1 }{ x^2+x+1 }-\frac{ 3 }{ x^2+x+1 } \right)\]\[\large =-\frac{ 1 }{ 3 }\frac{ 2x+1 }{ x^2+2x+1 }+\frac{ 1 }{ (x+1/2)^2+(\sqrt{3/4})^2 }\]

OpenStudy (anonymous):

wrong you failed at subdividing again you

OpenStudy (sirm3d):

uhm, that should be \[\large -\frac{ 1 }{ 3 }\frac{ 2x+1 }{ x^2+x+1 }+\frac{ 1 }{ (x+1/2)^2+(\sqrt{3/4})^2 }\]

OpenStudy (sirm3d):

the first term is the ln function, the second term is arctan function

OpenStudy (anonymous):

the answer is